OpenStudy (anonymous):

Solve the following equation: 4(5x - 9)square = 49

OpenStudy (anonymous):

(5x-9^2 = 49/4 taking square roots 5x-9 = +- 7/2 5x - 9 = 7/2 5x = 12.5 x = 2.5 or 5x-9 = -3.5 5x = 5.5 x = 1.1 solution set (1.1,2.5)

OpenStudy (anonymous):

Thanks!

OpenStudy (king):

yeah he is rite!!!!!!

OpenStudy (anonymous):

yw - ty king!!